You count 120 bacterial colonies. Dilutions: 1.0ml into 9ml, 0.1ml into 9.9mL and 0.5ml into 4.5 ml. You spread 0.2ml of your final dilution

Question: How many mls of the final dilution would you have needed to plate to get about 60 colonies?
Answer #1

dilution obtained by the first step=aliqout volume/final volume=1/10

dilution obtained at the second step= aliquot volume/final volume=0.1/10=1/100

dilution obtained at the third step=aliquot volume/final volume= 0.5/5=1/10 ( multiply with 2 on both neumerator and denominator)

total dilution= production of all individual dilutions

total dilution=1/10\times1/100\times1/10=1/10000

volume plated= 0.2 mL

CFU/mL= number of colonies/ ( volume plated \times dilution)

= 120/(0.2\times1/10000)

= 6\times10^6/ mL

to get 60 colonies

CFU/mL=number of colonies/ ( volume plated \times dilution)

volume plated= number of colonies/ (CFU/mL\timesdilution)

= 60/( 6\times10^6\times1/10000)

= 0.1 mL

volume of the final dilution would you have needed to plate to get about 60 colonies = 0.1 mL